### v1

# Distribution of lengths of closed saddle connections on moduli space of large genus translation surface

Abstract.
Let $S_{g}$ be a closed surface of genus $g$ and ${\mathcal{H}}_{g}$ be the moduli space of Abelian differentials on $S_{g}$. A stratum of ${\mathcal{H}}_{g}$, together with the Masur-Veech measure, becomes a probability space. Then the number of closed saddle connections with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ on a random flat surface in the stratum is a random variable. We prove that when $g\to\infty$, the distribution of the random variable converges to a Poisson distributed random variable. This result answers a question of Masur, Rafi and Randecker.
Keywords: Random flat surface; closed saddle connection
1. Introduction
In 2024 , Masur, Rafi and Randecker study the distribution of open saddle connections on a random translation surface in [11]. In this paper we study the distribution of closed saddle connections.
A translation surface can be denoted by $(X,\omega)$ where $X$ is a closed Riemann surface of genus $g$ and $\omega$ is an Abelian differential on $X$. $(X,\omega)$ has a flat structure defined by the conical metric $|\omega|$, with singularities at the zeros of $\omega$. If an $|\omega|$-geodesic of $(X,\omega)$ connects two zeros (or one zero to itself) of $\omega$ and has no other zeros in its interior, we call it an open (or closed) saddle connection.
Let ${\mathcal{H}}_{g}$ be the moduli space of unit-area translation surfaces of genus $g$. ${\mathcal{H}}_{g}$ has a natural stratification according to the zero order of Abelian differentials. Particularly, the stratum consisting of all abelian differentials in ${\mathcal{H}}_{g}$ which have only simple zeros is called principal stratum, denoted by ${\mathcal{H}}_{g}(1^{2g-2})$. For convenience, if a stratum consists of Abelian differentials on which the number of zeros of order $d_{i}$ is $\kappa_{i}$ we denote it by
${\mathcal{H}}_{g}(d_{1}^{\kappa_{1}},\cdots,d_{k}^{\kappa_{k}})$
where $\sum^{k}_{i=1}\kappa_{i}d_{i}=2g-2$.
Together with a finite natural measure which is introduced by Masur and Veech denoted by $\mu_{MV}$, every stratum ${\mathcal{H}}_{g}(\kappa)$ of ${\mathcal{H}}_{g}$ becomes a probability space with probability measure $\frac{\mu_{MV}}{\operatorname{Vol}{{\mathcal{H}}_{g}(\kappa)}}$, where $\operatorname{Vol}{{\mathcal{H}}_{g}(\kappa)}$ is the measure of total stratum, which we call the Masur-Veech volume of ${\mathcal{H}}_{g}(\kappa)$.
Given $(X,\omega)\in{\mathcal{H}}_{g}(1^{2g-2})$ and $a,b\in\mathbb{R}^{+}$, let $N_{g,[a,b]}(X,\omega)$ be the number of open saddle connections with lengths in the interval $\frac{a}{g},\frac{b}{g}]$ on $(X,\omega)$. Then $N_{g,[a,b]}:{\mathcal{H}}_{g}(1^{2g-2})\to\mathbb{N}_{0}$ becomes a random variable. In [11], Masur, Rafi and Randecker proved :
Theorem 1.1.
Let $a_{1},b_{1}],\cdots,[a_{k},b_{k}]$ be $k$ disjoint intervals. Then as $g\to\infty$, the vector of random variables
$(N_{g,[a_{1},b_{1}]},\cdots,N_{g,[a_{k},b_{k}]}):{\mathcal{H}}_{g}(1^{2g-2})\to\mathbb{N}_{0}^{k}$
converges jointly in distribution to a vector of random variables with Poisson distributions of means $\lambda_{[a_{i},b_{i}]}$, where
$\lambda_{[a_{i},b_{i}]}=8\pi(b_{i}^{2}-a_{i}^{2})$
for $i=1,\cdots,k$. That is,
$\mathbb{P}(N_{g,[a_{1},b_{1}]}=n_{1},\cdots,N_{g,[a_{k},b_{k}]}=n_{k})=\prod_{i=1}^{k}\frac{\lambda_{[a_{i},b_{i}]}^{n_{i}}e^{-\lambda_{[a_{i},b_{i}]}}}{n_{i}!}$
Their method is based on the relationship between Poisson distribution and its factorial moment [3], which was used first in the work of Mirzakhani–Petri [13] to study the distribution of closed hyperbolic geodesics on random hyperbolic surfaces in Teichmüller space with respect to Weil-Peterson measure.
As for translation surface, to use the method, [11] introduces a surgery to collapse the open saddle connections to obtain a new general translation surface in a new stratum. This surgery does not work for closed saddle connections and they proposed the distribution question in the situation of closed saddle connections, which inspires this work. This paper gives a new surgery to collapse a closed saddle connection, which ensures the same method can be used for closed saddle connections.
Let $L_{g,[a,b]}(X,\omega)$ be the number of closed saddle connections on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. Denote by $\mathcal{C}_{g}(\kappa)[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ the subset of ${\mathcal{H}}_{g}(\kappa)$ consisting of flat surfaces having closed saddle connections with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. The main result of this paper is the following theorem
Theorem 1.2.
For the stratum ${\mathcal{H}}(2^{g-1})$ and disjoint intervals $a_{1},b_{1}],\cdots,[a_{k},b_{k}]$, when $g\to\infty$, the random variable sequence
$(L_{g,[a_{1},b_{1}]},\cdots,L_{g,[a_{k},b_{k}]}):{\mathcal{H}}(2^{g-1})\to\mathbb{N}^{k}_{0}$
converges jointly in distribution to a vector of random variables with Poisson distributions of means $\lambda_{[a_{i},b_{i}]}$, where
$\lambda_{[a_{i},b_{i}]}=3\pi(b_{i}^{2}-a_{i}^{2}).$
Remark 1.3.
We will expalin the question is trivial for the principal stratum in Section 3: the probability measure of $\mathcal{C}_{g}(1^{2g-2})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ is zero when $g\to\infty$. For principal stratum, the length interval to be considered should be $[a,b]$.
Corollary 1.4.
Let ${\mathcal{H}}_{g}(2^{O(g)},1^{2g-2-2O(g)})$ be a stratum with $O(g)$ double zeros, then the result in Theorem 1.2 is also true.
2. Background
2.1. Moduli space of Abeilian differientials
Let $S_{g}$ be a closed surface of genus $g$, $\mathcal{M}_{g}$ be the moduli space, i.e, the set of isotopy classes of hyperbolic metric on $S_{g}$.
For some $X\in\mathcal{M}_{g}$, an Abelian differential $\omega$ on $X$ is a differential $1$-form which under a local coordinate $(U,z)$ of $X$ can be written by
$\omega=\omega(z)dz,$
where $\omega(z):U\to\mathbb{C}$ is a holomorphic function. The holomorphic differential space on $X$ is denoted by $H^{1}(X)$, which is a complex linear space of dimension $g$.
The moduli space of Abelian differential ${\mathcal{H}}_{g}$ on $S_{g}$ consists of $(X,\omega)$, where $X\in\mathcal{M}_{g}$ and $\omega\in H^{1}(X)$. ${\mathcal{H}}_{g}$ is a bundle over $\mathcal{M}_{g}$. Since $dimH^{1}(X)=g,dim\mathcal{M}_{g}=3g-3$, we have $dim{\mathcal{H}}_{g}=4g-3$. For $(X,\omega)$, define its area by $\int_{X}|\omega|^{2}$.
2.1.1. Flat metric
For $(X,\omega)\in{\mathcal{H}}_{g}$, it defines a conical metric with singularities on $X$ given by
$|\omega|=|\omega(z)||dz|.$
Consider the local coordinate given by $\omega$. Under the natural coordinate of $\omega$, the transition map is like
$z\to z+c$
for a constant $c$. We then call $(X,\omega)$ a translation surface and ${\mathcal{H}}_{g}$ becomes the moduli space of translation surfaces.
Denote the zero set of $(X,\omega)$ by $\Sigma$, in the natural coordinate, the flat metric becomes $|dz|$ on $X\setminus\Sigma$. Around a zero of order $n$, the total angle is $2(n+1)\pi$. Using polar doordinates the metric becomes
$ds^{2}=dr^{2}+(n+1)^{2}r^{2}d\theta^{2}.$
A saddle connection on $(X,\omega)$ is an $|\omega|$- geodesic segment connecting two zeros of $\omega$ and has no other zeros in the interior. We call a saddle connection open if it connects two different zeros and closed if it connects a zero to itself. Generally, a geodesic on $(X,\omega)$ is a concatenation of some saddle connections, with the angle between two adjacent saddle connections larger than $\pi$. However, there are geodesics with no zeros on them, which we call regular.
For a simple closed curve $\gamma$, there are two situations:
  • 1
    If there are regular geodesics homotopic to $\gamma$, then all such geodesics foliate a flat cylinder, with boundaries not regular.
  • 2
    There is a unique geodesic homotopic to $\gamma$.
For a saddle connection or a closed curve $\gamma$, denote its flat length on $(X,\omega)$ by $\ell_{\omega}(\gamma)$.
2.1.2. Stratum and Masur-Veech measure
For $(X,\omega)\in{\mathcal{H}}_{g}$, the sum of its zeros orders is 2g22g-2. ${\mathcal{H}}_{g}$ has a natural stratum structure by the orders of zeros: let $\kappa=(\kappa_{1},\cdots,\kappa_{n})$ be a set of $n$ positive integers satisfying $\sum_{i=1}^{n}\kappa_{i}=2g-2$. Let ${\mathcal{H}}_{g}(\kappa)$ be the subset of ${\mathcal{H}}_{g}$ consisting of Abelian differentials with zeros of orders $\kappa_{1},\cdots,\kappa_{n}$. Then we call ${\mathcal{H}}_{g}(\kappa)$ a stratum of ${\mathcal{H}}_{g}$ with characterization $\kappa$. When $\kappa=(1^{2g-2})$, we call ${\mathcal{H}}_{g}(\kappa)$ principal stratum. Then the total moduli space ${\mathcal{H}}_{g}$ can be decomposed by
${\mathcal{H}}_{g}=\bigcup_{\kappa}{\mathcal{H}}_{g}(\kappa),$
where $\kappa$ takes on all sets of positive integers with sum 2g22g-2.
Fix a characterization $\kappa=(\kappa_{1},\cdots,\kappa_{n})$, the local chart of ${\mathcal{H}}_{g}(\kappa)$ can be given by period mapping. For $(X,\omega)\in{\mathcal{H}}_{g}(\kappa)$, denote its zero set by $\Sigma$ and $\{\gamma_{1}\cdots\gamma_{d}\}$ be a basis of the relative homology $H_{1}(X,\Sigma,\mathbb{Z})$. Then the period mapping can be defined on a neighborhood of $(X,\omega)$ by
$(X^{\prime},\omega^{\prime})\to\left(\int_{\gamma_{i}}\omega^{\prime}\right)_{i=1,\cdots,d},$
where $(X^{\prime},\omega^{\prime})$ is in a neighborhood of $(X,\omega)$. Thus, locally ${\mathcal{H}}_{g}(\kappa)$ can be identified with $H_{1}(X,\Sigma,\mathbb{C})$.
From the dimension of $H_{1}(X,\Sigma,\mathbb{Z})$, we have
$dim_{\mathbb{C}}{\mathcal{H}}_{g}(\kappa)=dimH_{1}(X,\Sigma,\mathbb{C})=d=2g-1+n.$
So the dimension of principal stratum is equal to the total moduli space.
On each stratum, the period mapping induces a natural Lebesgue measure such that the lattice $H_{1}(X,\Sigma,\mathbb{Z}\oplus\mathbb{Z}i)$ has co-volume $1$, which is introduced by Masur [10] and Veech [16], we call it Masur-Veech measure. For a measured set $K$ of ${\mathcal{H}}_{g}(\kappa)$, denote its Masur-Veech measure by $\operatorname{Vol}(K)$.
Now we consider the hypersurface ${\mathcal{H}}^{1}_{g}(\kappa)$ of ${\mathcal{H}}_{g}(\kappa)$ consisting of $(X,\omega)$ with unit-area. The Masur-Veech measure on ${\mathcal{H}}^{1}_{g}(\kappa)$ can be defined as follows: for an open set $K$ on ${\mathcal{H}}^{1}_{g}(\kappa)$, its Masur-Veech measure is defined by
$\operatorname{Vol}(K)=dim_{\mathbb{R}}{\mathcal{H}}_{g}(\kappa)\cdot\operatorname{Vol}(Cone(K)),$
where $Cone(K)$ is the cone of $K$. In the following by abuse of notation we still denote ${\mathcal{H}}^{1}_{g}(\kappa)$ by ${\mathcal{H}}_{g}(\kappa)$.
Remark 2.1.
Note that the stratum ${\mathcal{H}}_{g}(\kappa)$ may not be connected, here we consider the connected component. For more details one can refer to [8].
Aggarwal [2, Theorem 1.4] proved that
Theorem 2.2.
For $g>1$ and a stratum ${\mathcal{H}}_{g}(\kappa)$ with characterization $\kappa=(\kappa_{1},\cdots,\kappa_{n})$,
$\frac{4}{\prod_{i=1}^{n}(\kappa_{i}+1)}\left(1-\frac{2^{2^{200}}}{g}\right)\leq\operatorname{Vol}({\mathcal{H}}_{g}(\kappa))\leq\frac{4}{\prod_{i=1}^{n}(\kappa_{i}+1)}\left(1+\frac{2^{2^{200}}}{g}\right).$
2.2. Siegel-Veech constant
2.2.1. Configuration
For a flat surface $(X,\omega)$ and a saddle connection $\gamma$ on it, define the holonomy vector of $\gamma$ to be $\int_{\gamma}\omega$ and denote it by $hol(\gamma)$. Denote the set of all saddle connections of $(X,\omega)$ by $\tilde{V}_{sc}(X,\omega)$. Then define
$V_{sc}(X,\omega)=\{hol(\gamma):\gamma\in\tilde{V}_{sc}(X,\omega)\}.$
In fact all holonomy vectors of $\gamma\in\tilde{V}_{sc}(X,\omega)$ define a holonomy mapping
$hol:\tilde{V}_{sc}(X,\omega)\to V_{sc}(X,\omega).$
By [18, Proposition 3.1], $V_{sc}(X,\omega)$ is a discrete subset of $\mathbb{R}^{2}$ for any $(X,\omega)$. Note that different saddle connections may have same holonomy vector. For example two homologous saddle connections must have same holonomy, here homologous for two open sadddle connections $\beta$ and $\gamma$ means $\beta\circ\gamma^{-1}$ is homologous to zero. For a vector $\vec{v}\in V_{sc}(X,\omega)$, define the multiplicity of $\vec{v}$ to be $|hol^{-1}(\vec{v})|$. It is known that almost any flat surface in any stratum does not have a pair of non-homologous sharing a same holonomy vector by [7, Proposition 3.1]. So in the following we consider only the surfaces on which if two saddle connections have same holonomy, then they are homologous.
To determine how many surfaces have some specific saddle connections, we need to classify the saddle connections by some geometric data, which we call configuration.
For open saddle connections, first fix two zeros $p_{1}$ and $p_{2}$ with orders $m_{1}$ and $m_{2}$, consider the saddle connections connecting $p_{1}$ and $p_{2}$ of multiplicity $p$ and denote them by $\gamma_{1},\cdots,\gamma_{p}$, where the orientation is from $p_{1}$ to $p_{2}$ and the cyclic order of $\gamma_{i}$ is clockwise at $p_{1}$. Now we consider the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{1}$ and $p_{2}$, which are even multiples of $\pi$. Let the the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{1}$ be $2(a^{\prime}_{i}+1)\pi$ , the the angle between $\gamma_{i}$ and $\gamma_{i+1}$ at $p_{2}$ be $2(a^{\prime\prime}_{i}+1)\pi$. Then a configuration $\mathcal{C}$ of open saddle connections consists of $(m_{1},m_{2},a^{\prime}_{i},a^{\prime\prime}_{i})$.
The situation of closed saddle connections is more complicated. Again we consider the closed saddle connections of multiplicity $p$. Suppose $\gamma_{1},\cdots,\gamma_{p}$ are homologous, and $\gamma_{1}$ connects $p_{1}$ to itself. Note that some $\gamma_{i}$ may connect other zero to itself than $p_{1}$, and some closed saddle connections may bound cylinders. Denote by $J=\{j_{1},\cdots,j_{l}\}\in\{1,2,\cdots,p\}$ the index set such that $\gamma_{j_{1}},\cdots,\gamma_{j_{l}}$ bound $q$ cylinders. Cut off the $q$ cylinders and $\gamma_{1},\cdots,\gamma_{p}$, there remain $p$ connected components $X_{1},\cdots,X_{p}$. $X_{1},\cdots,X_{p}$ can be divided into two types: components of type one have two boundary components, type two have one, which is constructed by two $\gamma_{i}$ and $\gamma_{j}$ sharing one zero (and don’t be cut off). Consider the angles the boundary curves bound inside the components. If some $X_{k}$ is of type one with boundaries $\gamma_{k}$ and $\gamma_{k+1}$, denote the angles by $(2b^{\prime}_{k}+1)\pi$ and $(2b^{\prime\prime}_{k}+1)\pi$. If some $X_{i}$ is of type two with boundary constructed by $\gamma_{i}$ and $\gamma_{i+1}$, denote the angles by $2(a^{\prime}_{i}+1)\pi$ and $2(a^{\prime\prime}_{i}+1)\pi$. Then a configuration $\mathcal{C}$ of closed saddle connections consists of $(J,b^{\prime}_{k},b^{\prime\prime}_{k},a^{\prime}_{i},a^{\prime\prime}_{i})$.
For a configuration $\mathcal{C}$, denote by $V_{\mathcal{C}}(X,\omega)$ be the set of holonomy vectors of saddle connections satisfying $\mathcal{C}$.
2.2.2. Siegel-Veech formula and Siegel-Veech constant
Given a stratum ${\mathcal{H}}_{g}(\kappa)$ and a configuration $\mathcal{C}$. Let $f:\mathbb{R}^{2}\to\mathbb{R}$ be a integrable function with compact support, the Siegel-Veech transform can be defined by
$\hat{f}_{\mathcal{C}}$ $:{\mathcal{H}}_{g}(\kappa)\to\mathbb{R}$
$\hat{f}_{\mathcal{C}}(X,\omega)$ $=\sum_{\vec{v}\in V_{\mathcal{C}}(X,\omega)}f(\vec{v}).$
Theorem 2.3 (Siegel-Veech formula [17]).
Let ${\mathcal{H}}_{g}(\kappa)$ be a connected stratum and $\mathcal{C}$ a configuration of saddle connections. Then there exists a constant $c(\mathcal{C},{\mathcal{H}}_{g}(\kappa))$ such that for every integrable $f:\mathbb{R}^{2}\to\mathbb{R}$ with compact support,
$\mathbb{E}(\hat{f}_{\mathcal{C}})=\frac{1}{\operatorname{Vol}({\mathcal{H}}_{g}(\kappa))}\int_{{\mathcal{H}}_{g}(\kappa)}\hat{f}_{\mathcal{C}}d\mu_{MV}=c(\mathcal{C},{\mathcal{H}}_{g}(\kappa))\int_{\mathbb{R}^{2}}fdxdy.$
The constant $c(\mathcal{C},{\mathcal{H}}_{g}(\kappa))$ is called Siegel-Veech constant of $\mathcal{C}$ and ${\mathcal{H}}_{g}(\kappa)$.
There have been many asymptotics of Siegel–Veech constants for large genus, one can see [5, 14, 1, 2, 4] for more results. Here we only state a result about closed saddle connections, for more details refer to [15].
Theorem 2.4 ([15] Appendix Poposition A.1).
Let ${\mathcal{H}}_{g}(\kappa)$ be a connected stratum and a fixed zero $\sigma$ of order $m$. For any configuration $\mathcal{C}$ of closed saddle connections at $\sigma$ of multiplicity $p$ involving $q$ cylinders and $t\leq p-1$ distinct singularities $\sigma_{1},\cdots,\sigma_{t}$ of orders $m_{1},\cdots,m_{t}$ other than $\sigma$, the Siegel-Veech constant associated to $\mathcal{C}$ is
$c(\mathcal{C},{\mathcal{H}}_{g}(\kappa))=\frac{(m+1)\prod_{i=1}^{t}(m_{i}+1)}{2^{p-1}|\Gamma_{-}||\Gamma|(p+q-1)!}(1+o(1))^{p}O(\frac{1}{g^{2p+q-2}}),$
where $|\Gamma_{-}|$ and $|\Gamma|$ are combinatorial constants depending on $p$.
2.3. Factorial moment
For a probability space $(\Omega,\mathbb{P})$ and a random variable $X$ on it, its factorial moment generating function (FMGF) is defined as:
$M_{X}(t)=\mathbb{E}[(1+t)^{X}],$
and its k-th factorial moment $\mathbb{E}[(X)_{k}]$ is defined by
$\mathbb{E}[(X)_{k}]=\mathbb{E}[X(X-1)\cdots(X-k+1)].$
In fact it is the k-th derivative of $M_{X}(t)$ with respect to $t$ at $t=0$.
Recall that a random variable $X:\Omega\to\mathbb{N}_{0}$ is Poisson distributed with mean $\lambda$ if
$\mathbb{P}(X=k)=\frac{\lambda^{k}e^{\lambda}}{k!}.$
The following theorem from [3] gives the relationship between factorial moment and distribution for Poisson distribution:
Theorem 2.5 (The method of moment from [3], Theorem 1.23).
Let $\left\{(\Omega_{i},\mathbb{P}_{i})\right\}_{i\in\mathbb{N}}$ be a sequence of probability spaces. For $m\in\mathbb{N}$, let $N_{1,i},\cdots,N_{m,i}:\Omega_{i}\to\mathbb{N}_{0}$ be random variables for all $i\in\mathbb{N}$ and suppose there exist $\lambda_{1},\cdots,\lambda_{m}\in(0,\infty)$ such that:
$\lim_{i\to\infty}\mathbb{E}[(N_{1,i})_{k_{1}}\cdots(N_{m,i})_{k_{m}}]=\lambda_{1}^{k_{1}}\cdots\lambda_{m}^{k_{m}}$
for all $k_{1}\cdots k_{m}\in\mathbb{N}$. Then
$\lim_{i\to\infty}\mathbb{P}[(N_{1,i})=n_{1},\cdots,(N_{m,i})=n_{m}]=\prod^{k}_{j=1}\frac{\lambda_{j}^{n_{j}}e^{-\lambda_{j}}}{n_{j}!}.$
That is, $(N_{1,i},\cdots,N_{m,i}):\Omega_{i}\to\mathbb{N}^{m}_{0}$ converges jointly in distribution to a vector which is independently Poisson distributed with means $\lambda_{1},\cdots,\lambda_{m}$.
3. Some claims
Before the proof of the main theorem, we have to make some claim which will simplify the situation. We consider the stratum whose zero order is $O(1)$, and it inplies $|\kappa|=O(g)$.
Claim 1.
A closed saddle connection is non-separable.
Proof.
For a closed connection $\gamma$ of $(X,\omega)$, if it’s separable, then it cut $X$ into $X_{1}$ and $X_{2}$, and $\partial X_{1}=\partial X_{2}=\gamma$. By Stokes theorem, $\int_{\gamma}\omega=0$, but since $\gamma$ is a saddle connection its length is $|\int_{\gamma}\omega|$, which is a contradiction. ∎
Claim 2.
The probability measure of the subset of ${\mathcal{H}}_{g}(\kappa)$ which has a closed saddle connection involving $q$ cylinders of multiplicity $p$ goes to 0 as $g\to\infty$, where $p>1$ or $q>0$.
Proof.
From Theorem 2.4, the Siegel-Veech constant of configuration with closed saddle connection involving $q$ cylinders of multiplicity $p$ is $O(\frac{1}{g^{2p+q-2}})$ . Denote by $\mathcal{C}^{p,q}_{g}(\kappa)[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ the subset of ${\mathcal{H}}_{g}(\kappa)$ which has a closed saddle connection of length in the interval $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ involving $q$ cylinders of multiplicity $p$. Then for $p\geq 2$ or $q\geq 1$, we have
$\operatorname{Vol}(\mathcal{C}^{p,q}_{g}(\kappa)[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}])\leq|\kappa|O(\frac{1}{g^{q}})\frac{b^{2}-a^{2}}{g}\operatorname{Vol}({\mathcal{H}}_{g}(\kappa))=O(\frac{1}{g})\operatorname{Vol}({\mathcal{H}}_{g}(\kappa)$
where the coefficient $|\kappa|$ is the selection of the zero that the closed saddle connection goes through. So when $g\to\infty$ its probability measure goes to 0 and it suffices to consider the closed saddle connection with multiplicity $1$ and no cylinders around it. ∎
Claim 3.
The angle of a closed saddle connection at the zero which it connects is odd multiples of $\pi$.
Proof.
For an Abelian differential $(X,\omega)$, if it has a closed saddle connection $\gamma$ connecting a zero $p$, if its angle at $p$ is $2k\pi$, then the holonomy of $\gamma$ and $\gamma$ have same direction. But
$\int_{\gamma}\omega=-\int_{-\gamma}\omega.$
It implies $\int_{\gamma}\omega=0$, which is a contradiction. ∎
Claim 4.
If a closed saddle connection has angle $\pi$ at one side, it must have a cylinder at the side.
Proof.
Let $\gamma$ be a colsed saddle connection on $(X,\omega)$ which has angle $\pi$ at one side. Since $(X,\omega)$ is oriented, its normal bundle of $\gamma$ at the side is oriented. Consider the exponential map from the bundle to $(X,\omega)$, which is well-defined on $\gamma\times[0,s]$ for some $s$ small sufficiently and the image is a cylinder.
Combine claim $2,3$ and $4$, we have that for a closed saddle connection $\gamma$ on $(X,\omega)$ in principal stratum, since its total angle around the zero is $4\pi$ and has to be divided into odd multiples of $\pi$, it must have angle $\pi$ at one side of $\gamma$ and $3\pi$ at another. So there is a cyliner with $\gamma$ as a boundary. As for the other boundary, if it consists of some open saddle connections, then these saddle connections have angles $\pi$. That means there are at least two non-homologous saddle connections have angle $\pi$. The measure of the set of such surfaces is zero. So we only need to consider the cylinder bounded by curves homologous to $\gamma$. Then from Claim 2 the probability measure of $\mathcal{C}_{g}(1^{2g-2})$ has limit
$\lim_{g\to\infty}\frac{\operatorname{Vol}(\mathcal{C}_{g}(1^{2g-2})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}])}{\operatorname{Vol}({\mathcal{H}}_{g}(1^{2g-2})}=O(\frac{1}{g})\to 0.$
So we need consider the question on stratum with higher order zeros. Next we will consider ${\mathcal{H}}_{g}(2^{g-1})$ firstly.
4. Surgery
This section we will introduce some surgeries that, when combined, can collapse a closed saddle connection.
4.1. Open up a higher-order zero
Masur, Rafi and Randecker introduce a surgery in [11] which is a variation of a surgery from [7]. This surgery can collapse a saddle connection which is not closed and has multiplicity $1$.
Let $(X,\omega)$ be an Abelian differential, $\sigma$ be an open saddle connection with endpoints $v_{1}$ and $v_{2}$, whose orders are $n_{1}$ and $n_{2}$ respectively. Since the total angle at $v_{1}$ is $2(n_{1}+1)\pi$, one can extend $\sigma$ from $v_{1}$ along its direction to $v_{3}$ and denote by $\sigma^{\prime}$ the geodesic segment $v_{1}v_{3}$ such that $\ell_{\omega}(\sigma)=\ell_{\omega}(\sigma^{\prime})$ and the angle between $\sigma$ and $\sigma^{\prime}$ is $2k_{1}\pi$ and $2k_{2}\pi$, where $k_{1}+k_{2}=n_{1}+1$. If $\sigma^{\prime}$ does not go through other zeros of $(X,\omega)$, which means $\sigma^{\prime}$ is a ray from $v_{1}$, the following surgery can be carried out.
Cut along $\sigma+\sigma^{\prime}$ and denote the two copies of $\sigma+\sigma^{\prime}$ and $v_{1}$ by $\sigma^{\pm}+(\sigma^{\prime})^{\pm}$ and $v_{1}^{\pm}$. Then glue $\sigma^{+}$ and $(\sigma^{\prime})^{+}$, $\sigma^{-}$ and $(\sigma^{\prime})^{-}$. This surgery reduces the order of $v_{1}$ and constructs a new Abelian differential $(X^{\prime},\omega^{\prime})$. On $(X^{\prime},\omega^{\prime})$, the total angle at $v_{1}^{+}$, $v_{1}^{-}$ and $v_{2}$ are $2k_{1}\pi$, $2k_{2}\pi$, and $2(n_{2}+2)\pi$ respectively, that is, the order of $v_{1}^{+}$, $v_{1}^{-}$ and $v_{2}$ are $k_{1}-1$, $k_{2}-1$, and $n_{2}+1$. Particularly, if $v_{1}$ is a simple zero, the surgery collapses the saddle connection. We call this surgery collapsing surgery.
Figure 1. The collapsing surgery
In our situation, we need to use the inverse of collapsing surgery which we call it opening surgery. From Claim 2, it suffices to consider the subset $\mathcal{C}^{1,0}_{g}(2^{g-1})$. For an Abelian differential $(X,\omega)$ in $\mathcal{C}^{1,0}_{g}(2^{g-1})$, choose a closed saddle connection $\gamma$ at a double zero $p$. We can reverse the surgery introduced above as follows.
Since the total angle at $p$ is $6\pi$, from Claim 3 and Claim 4, the angles on both side of $\gamma$ at $p$ must be $3\pi$. Fix the orientation of $\gamma$, choose two rays $\sigma^{+}$ and $\sigma^{-}$ from $p$ such that $\ell_{\omega}(\sigma^{+})=\ell_{\omega}(\sigma^{-})=\ell_{\omega}(\gamma)$, and the angle between $\gamma$ and $\sigma^{+}$ is $\pi$, the angle between $\gamma$ and $\sigma^{-}$ is $\pi$, where the sign is consistent with the orientation of the surface. Denote by $q^{+}$ and $q^{-}$ the other endpoints of $\sigma^{+}$ and $\sigma^{-}$. Then cut along $\sigma^{+}+\sigma^{-}$ and glue the two copies of $\sigma^{+}$ and $\sigma^{-}$. After the surgery, the two copies of $p$ become a regular and a simple zero, $q^{+}$ and $q^{-}$ are glued to become a new simple zero $q$.
Figure 2. The opening surgery
This surgery is the inverse of the collapsing surgery above : the double zero is replaced by two simple zeros, and it can be carried out except for the surgery locus goes through some zero, that is, it has two non-homologous saddle connections with angle $2\pi$. From period mapping, such subset has measure zero. We call the subset on which the surgery can be carried out permissible set, and denote it by $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$, from the discussion above we have
Proposition 4.1.
$\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$ is a full measure subset of $\mathcal{C}^{1,0}_{g}(2^{g-1})$.
By the surgery of collapsing this map is one-to-one and denote the opening surgery by $F_{1}$.
Figure 3. The resulting surface
Remark 4.2.
We can also choose $\sigma^{\pm}$ to make the angle between $\gamma$ and them are $2\pi$, it suffices to make sure the opening operation can obtain an abelian differential and a smooth loop homotopic to $\gamma$.
4.2. Move zero along closed curve and pinch
We have constructed the mapping $F_{1}$, which takes a closed saddle connection $\gamma$ at a double zero to two saddle connections $\gamma_{1}$ and $\gamma_{2}$ sharing same endpoints $p$ and $q$ which are both simple zeros. Moreover, the angles between the two saddle connections at the two simple zeros (on both sides) are all $2\pi$, and if we denote by $\ell_{\omega}(\gamma)=L$, we have $\ell_{\omega^{\prime}}(\gamma_{1})=L$, $\ell_{\omega^{\prime}}(\gamma_{2})=2L$, where $(X^{\prime},\omega^{\prime})=F_{1}(X,\omega)$. If we want to collapse $\gamma_{1}$ and $\gamma_{2}$ simultaneously, we have to make a surgery to move $q$ along $\gamma_{2}$ to adjust the length of $\gamma_{2}$.
4.2.1. Move the zero locally by period mapping.
First we need to choose a special basis of $H^{1}(X,\Sigma;\mathbb{C})$, where $\Sigma=\{p,q,p_{2},\cdots,p_{g-1}\}$ is the zero set of $(X^{\prime},\omega^{\prime})=F_{1}(X,\omega)$. From Claim 1, the homology class $\gamma]=[\gamma_{1}+\gamma_{2}]$ is non-separable. So we can choose a basis of $H^{1}(X,\mathbb{C})$
$(\alpha_{1},\beta_{1},\cdots,\alpha_{g},\beta_{g}),$
where $\alpha_{1}=[\gamma]$.
Next we choose the relative homology class
$(pq,pp_{2}\cdots,pp_{g-1}),$
where $\gamma_{1}]=pq$, and $pp_{i}$ is freely homotopic to 0 in $H^{1}(X,\mathbb{C})$. Together they compose a basis of $H^{1}(X,\Sigma;\mathbb{C})$, and their holonomy gives a period mapping around $(X^{\prime},\omega^{\prime})$ locally.
Let $\omega^{\prime}(\alpha_{1})=(x_{1},y_{1})$, $\omega^{\prime}(\beta_{1})=(x_{2},y_{2})$, $\omega^{\prime}(\alpha_{2})=(x_{3},y_{3})$ $\omega^{\prime}(\beta_{2})=(x_{4},y_{4})$. Now choose a relative cohomology class $\upsilon\in H^{1}(X,\Sigma;\mathbb{C})$ , which can be considered as a tangent vector in $T_{{\mathcal{H}}_{g}(1,1,\cdots,2)}(X^{\prime},\omega^{\prime})$ such that
(4.1) $\upsilon(\alpha_{1})=(-x_{1},-y_{1}),\upsilon(\alpha_{2})=(-\frac{y_{2}x_{1}}{y_{4}},\frac{x_{2}y_{1}}{x_{4}})$
and on the other basis we assign zero to $\upsilon$.
Remark 4.3.
Here $\upsilon(\alpha_{1})$ is chosen to guarantee that along the curve $(X^{\prime},\omega^{\prime})+t\upsilon$ in moduli space the zero moves along $\gamma$, and $\upsilon(\alpha_{2})$ is chosen to guarantee the resulting surface has area $1$, since the area of $(X,\omega)$ can be written by
$\int_{X}|\omega|^{2}=\frac{i}{2}\sum_{i}(\int_{\alpha_{i}}\omega\int_{\beta_{i}}\overline{\omega}-\int_{\beta_{i}}\omega\int_{\alpha_{i}}\overline{\omega}).$
Consider the curve in ${\mathcal{H}}_{g}(1,1,2,,\cdots,2)$: $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. If $\gamma_{1}$ and $\gamma_{2}$ don’t degenerate along the curve, then we obtain
$F_{2}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})\to{\mathcal{H}}_{g}(1^{2},2^{g-1})$
such that $F_{2}[(X,\omega)]=F_{1}[(X,\omega)]+\upsilon$. From the construction above the resulting surface has two saddle connections in the relative homology class $\gamma_{1}$ and $\gamma_{2}$ with equal length.
4.2.2. $F_{2}$ is well-defined almost everywhere
We have defined a mapping if the saddle connections $\gamma_{1}$ and $\gamma_{2}$ are preserved along the curve $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$ on $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$. Next we will see the mapping can be defined for almost every flat surface in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$.
From above we know the moving surgery can be realized except for some of $\gamma_{1}$ or $\gamma_{2}$ degenerate along $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,1]$. Suppose $T$ is the first time when the geodesic in $\gamma_{1}]$ can be represented by $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$, since the saddle connection is smooth on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$, the corner of $pp_{i}-p_{i}p_{j}-\cdots-p_{k}q$ must be $\pi$. But the holonomy of $pp_{i}$ is not changed on $(X^{\prime},\omega^{\prime})+t\upsilon,t\in[0,T]$ This implies there exists some $pp_{i}$ such that its holonomy on $(X^{\prime},\omega^{\prime})$ has the same direction with $p_{0}p_{1}$. And by the relation of holonomy under $F_{1}$, this implies $p_{0}p_{i}$ has the same direction with the closed curve on $(X,\omega)$. Under period mapping, on each local chart this is a measure-zero set. So all such $(X,\omega)$ is a subset of measure zero in ${\mathcal{H}}_{g}(2^{g-1})$, also in $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$.
4.2.3. Pinching
On the subset of $\mathcal{C}^{1,0}_{g,Per}(2^{g-1})$ where $F_{2}$ is well-defined, we can cut along $\gamma_{1}+\gamma_{2}$ and denote the two copies of $p$ and $q$ by $^{+}, qq^{-}$, $^{+}, pp^{-}$, the two copies of $\gamma_{1}$ and $\gamma_{2}$ by $\gamma^{+}_{1}$, $\gamma^{-}_{1}$, $\gamma^{+}_{2}$, $\gamma^{-}_{2}$. Notice that the total angles at $^{+}, qq^{-}$, $^{+}, pp^{-}$ are all $2\pi$, which means $\gamma^{\pm}_{1}$ have same direction with $\gamma^{\pm}_{2}$, then glue $\gamma^{+}_{1}$ with $\gamma^{+}_{2}$, $\gamma^{-}_{1}$ with $\gamma^{-}_{2}$, which make $^{+}, qq^{-}$, $^{+}, pp^{-}$ become regular points. Since $\gamma_{1}+\gamma_{2}$ is non-separable, this surgery reduces the genus by one, and obtains a new abelian differential on which we mark two points $p^{+}$ and $p^{-}$ so as to find its inverse, then we obtain:
$F_{3}:\mathcal{C}^{1,0}_{g,Per}(2^{g-1})\to{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$
Figure 4. Moving and cut
Figure 5. Pinching
4.2.4. $F_{3}$ is measure-preserving.
This surgery can be inverse except for some directions which has saddle connections, which is a measure zero subset $A_{0}(X,\omega)$ in $A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$ for every surface in ${\mathcal{H}}_{g-1}(2^{g-2},0,0)$: choose a vector $\kappa$ in $A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$, consider the two rays from $p^{+}$ and $p^{-}$ whose holonomy is $\kappa$, if the rays exist, then cut the surface along the two rays to obtain two closed loci and glue them. Let
$\tilde{{\mathcal{H}}}=\{(X,\omega,\tau):(X,\omega)\in{\mathcal{H}}_{g-1}(2^{g-2},0,0),\tau\in A_{0}(X,\omega)\},$
which is a measure zero subset of ${\mathcal{H}}_{g-1}(2^{g-2},0,0)\times A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]}$.
Moreover, under local chart determined by the basis we choose,
$F_{3}(z_{1},\cdots,z_{n})=(z_{1},z_{i}+\upsilon(z_{i})),$
whose Jacobian has deteminant one and thus is measure-preserving. So we have
$F_{3}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1}))=[{\mathcal{H}}_{g-1}(2^{g-2},0,0)\times(A_{[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]})]\setminus\tilde{{\mathcal{H}}}$,
Since $\operatorname{Vol}({\mathcal{H}}_{g-1}(2^{g-2},0,0))=\operatorname{Vol}({\mathcal{H}}_{g-1}(2^{g-2}))$, we have
Proposition 4.4.
$\operatorname{Vol}(\mathcal{C}^{1,0}_{g}(2^{g-1})))=\operatorname{Vol}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1})))=\frac{\pi(b^{2}-a^{2})}{g}\operatorname{Vol}({\mathcal{H}}_{g-1}(2^{g-2})).$
5. Collapse closed saddle connections simultaneously
From the construction above we can collapse one closed saddle connection on some $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$ and obtain a new surface $(X^{\prime},\omega^{\prime})\in{\mathcal{H}}_{g-1}(2^{g-2},0,0)$. If we want to collapse any $k$ closed saddle connections, we have to ensure there aren’t two short closed saddle connections sharing one zero. We need the following collar lemma [12, 6, 9].
5.1. Collar lemma
This section we will introduce the collar lemma on flat surface, and use it to prove the closed saddle connections we want to collapse are disjoint.
Definition 5.1.
Let $A$ be an annulus on $(X,\omega)$ such that there are no zeros of $\omega$ in the interior. We say that $A$ is expanding if the curvature of each boundary component of $A$ has constant sign, either positive or negative at each point, and the boundary curves are equidistant.
Denote the boundary components by $\partial_{0}A$ and $\partial_{1}A$. If $A$ is an expanding annulus, then we assume that the total curvature $\kappa(\partial A_{0})=\kappa(A)>0$ and $\kappa(\partial A_{1})=-\kappa(A)<0$. Let $d(A)$ be the distance between $\partial_{0}A$ and $\partial_{1}A$.
For $\gamma$ a simple closed curve , let $C(\gamma)$ be the maximal flat cylinder on $(X,\omega)$ with core curves homotopic to $\gamma$, and let $A(\gamma)$ be the maximal expanding annulus, that is, one of whose boundary components is a $\omega$-geodesic homotopic to $\gamma$.
Lemma 5.2.
[9, Lemma 1] Assume that $\gamma$ is a simple closed curve on $(X,\omega)$ such that its hyperbolic length $\ell_{X}(\gamma)$ is sufficiently small. Then
  1. (1)
    $\frac{1}{\ell_{X}(\gamma)}\asymp\max\{\operatorname{Mod}(A(\gamma)),\operatorname{Mod}(C(\gamma))\};$
  2. (2)
    $\operatorname{Mod}(A(\gamma))\asymp\log\frac{d(A(\gamma))}{\ell_{\omega}(\gamma)}.$
This lemma originates from Minsky [12, §4.3], moreover from the proof of Minsky we know in fact the comparability coefficient depends on the total curvature $\kappa(A(\gamma))$.
For simplicity, denote by $A=A(\gamma)$ and $C=C(\gamma)$. Either $\frac{1}{\ell_{X}(\gamma)}\asymp\operatorname{Mod}(C)$ or $\frac{1}{\ell_{X}(\gamma)}\asymp\operatorname{Mod}(A)$. In the first case, we have $\operatorname{Mod}(C)=\frac{d(C)}{\ell_{\omega}(\gamma)}$, where $d(C)$ denotes the width of $C$. Thus we have
(5.1) $d(C)\asymp\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)}.$
In the second case, we have
(5.2) $\log\frac{d(A)}{\ell_{\omega}(\gamma)}\asymp\operatorname{Mod}(A)\asymp\frac{1}{\ell_{X}(\gamma)},$
which implies
$d(A)\succ\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)}.$
We conclude that there is an annulus homotopic to $\gamma$ such that if $\ell_{X}(\gamma)$ is sufficiently small then the width of the annulus is at least
$c\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)},$
where $c>0$ is a constant depending on $\kappa(A(\gamma))$.
Proposition 5.3.
For $g$ large enough, let $\gamma_{1},\cdots,\gamma_{k}$ be $k$ closed saddle connections of multiplicity $1$ with no cylinder around them on $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$. If their length are all in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, $\gamma_{1},\cdots,\gamma_{k}$ are disjoint.
Proof.
By assumption, any $\gamma_{i}$ has only one geodesic representation. Since the minimal intersection between two homotopic class is realized by its geodesic representation, if $\gamma_{i}$ and $\gamma_{j}$ intersect, they are intersected essentially. From the collar lemma 5.2, $\gamma_{i}$ must go through the collar (in fact, it’s expanding) $A_{j}$ around $\gamma_{j}$. As $\gamma_{i}$ and $\gamma_{j}$ have only one double zero on them, their total curvatures have a uniform bound. Then the width of $A_{j}$ has a lower bound $c\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)}$, where $c$ is a constant. When $g$ goes to infinity, $\ell_{\omega}(\gamma_{i})$ goes to zero, and $\ell_{X}(\gamma_{i})$ must go to zero. So for $g$ large enough, we have
$c\frac{\ell_{\omega}(\gamma_{i})}{\ell_{X}(\gamma)}\geq c\frac{a}{\sqrt{g}\ell_{X}(\gamma)}\geq\frac{b}{\sqrt{g}},$
and we have $\ell_{\omega}(\gamma_{j})\geq\frac{b}{\sqrt{g}}$, which is a contradiction. ∎
5.2. Collapsing simultaneously
Now fix a positive integer $K$ and $n_{1},\cdots,n_{k}$ a partition of $K$. We want to compute the limit of the expectation
$(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}},$
for which we will collapse $K$ closed saddle connections in order.
Define $\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1}))=(X,\omega,\Gamma_{1},\cdots,\Gamma_{k})$, where $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$, and $\Gamma_{i}$ is an ordered list of $n_{i}$ closed saddle connections $(\gamma_{1,i},\cdots,\gamma_{n_{i},i})$ with lengths in $\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]$. Suppose $a_{1}\leq b_{1}\leq a_{2}\leq b_{2}\cdots\leq a_{k}\leq b_{k}$. Note that when choose $a=a_{1},b=b_{k}$, by Proposition 5.3, these closed saddle connections are disjoint when $g\to\infty$.
We will adjust the surgery from last section to collapse the closed saddle connections in $\Gamma_{1},\cdots,\Gamma_{k}$ in order. The opening and pinching operations are same to the situation of one curve. The only modification is the equation 4.1: by remark 4.3, we need to choose the moving vector $\upsilon$ to ensure the area unchanged, hence we can choose a basis containing $\gamma^{\prime}_{1},\gamma_{1}\cdots,\gamma^{\prime}_{K},\gamma_{K}$, where $\gamma^{\prime}_{i}$ is a saddle connection from opening $\gamma_{i}$. Then the moving vector $\upsilon$ can be constructed like equation 4.1: after defining the value on $\gamma_{i}$, we can choose another relative homotopy class and give a value of $\upsilon$ on it to ensure the area unchanged. This means we open and move the closed saddle connections simultaneously and then pinch them.
Finally we obtain a mapping
$\hat{F}:\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1}))\to M_{K}\cdot{\mathcal{H}}_{g-K}(2^{g-1-K})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}}$
where $M_{K}$ is a combinatorial data consisting of choosing $K$ ordered zeros from $g$ zeros to label. As the discussion above, the mapping is one-to-one and measure-preserving. So we have
$\int_{{\mathcal{H}}_{g}(2^{g-1})}(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}d\mu_{MV}$ $=\operatorname{Vol}(\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1})))$
$=|M_{K}|\prod^{k}_{i=1}[\frac{\pi(b^{2}-a^{2})}{g}]^{n_{i}}\operatorname{Vol}({\mathcal{H}}_{g-k}(2^{g-1-K}))$
$\to\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{g-1-K})),g\to\infty,$
where the limit is because $|M_{K}|=g\cdots(g-K+1)$ and
$\lim_{g\to\infty}\frac{g\cdots(g-K+1)}{g^{n_{1}+\cdots+n_{k}}}=\frac{g\cdots(g-K+1)}{g^{K}}=1.$
Then the limit of expectation is
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{g-1-K}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}$
From Theorem 2.2 ,
$\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{g-1-K}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}=3^{K}.$
So we have
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\prod^{k}_{i=1}[3\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}.$
By theorem 2.5, we get theorem 1.2.
From the proof above we can prove Corollary 1.4.
proof of Corollary 1.4.
Since we don’t need to deal with simple zeros in the collapsing surgery, samilarly to the proof of theorem 1.2, we have
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{O(g)-K},1^{2g-2-2O(g)}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{O(g)},1^{2g-2-2O(g)}))}\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}},$
and
$\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{g-1},1^{2g-2}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1},1^{2g-2}))}=3^{K}.$
Then
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\prod^{k}_{i=1}[3\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}.$
6. General stratum
This section we consider the stratum ${\mathcal{H}}_{g}(m^{O(g)},1^{2g-2-mO(g)})$, where $m\geq 3$.
If a stratum has zeros of order more than $3$, the angles at both sides are not certain. But we can consider some fixed configuration of closed saddle connection. For a configuration $(J,b^{\prime}_{k},b^{\prime\prime}_{k},a^{\prime}_{i},a^{\prime\prime}_{i})$ of closed saddle connection. We have known it suffices to consider multiplicity $1$ without cylinders, then the configuration become $(b^{\prime},b^{\prime\prime})$, where $b^{\prime}+b^{\prime\prime}=m$, which means the angles of the closed saddle connection are $(2b^{\prime}+1)\pi$ on one side and $(2b^{\prime\prime}+1)\pi$ on the other side. Now we choose all such configuration at each $m$-order zeros.
For $k$ such closed saddle connections, we can use the surgery in Section 4 to collapse them. Note that in this situation the opening surgery will replace the zero of order $m$ to two zeros of order $1$ and m1m-1, and the angles at both sides of the zero of m1m-1 order are $(2b^{\prime})\pi$ and $(2b^{\prime\prime})\pi$. Then after moving and pinching, the simple zero become two regular points and the zero of m1m-1 order become two zeros of order $b^{\prime}-1$ and $b^{\prime\prime}-1$. Finally we obtain a mapping
$\hat{F}:\hat{\mathcal{C}}^{1,0}_{g,Per}(m^{O(g)},1^{2g-2-mO(g)})\to{\mathcal{H}}_{g-K}((b^{\prime}-1)^{K},(b^{\prime\prime}-1)^{K},m^{O(g)-K},1^{2g-2-mO(g)})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}}.$
So we have
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\prod^{k}_{i=1}(\frac{(m+1)\pi(b_{i}^{2}-a_{i}^{2})}{b^{\prime}b^{\prime\prime}})^{n_{i}}.$
And by Theorem 2.5 we have
Theorem 6.1.
For the stratum ${\mathcal{H}}(m^{O(g)},1^{2g-2-mO(g)})$, a configuration $(b^{\prime},b^{\prime\prime})$ and $a,b\in\mathbb{R}$, let the random variable
$L_{\mathcal{C},g,[a,b]}:{\mathcal{H}}(m^{O(g)},1^{2g-2-mO(g)})\to\mathbb{N}_{0}$
be the number of closed saddle connections on $(X,\omega)$ satisfying the configuration with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. Then for disjoint intervals $a_{1},b_{1}],\cdots,[a_{k},b_{k}]$, when $g\to\infty$, the random variable sequence
$(L_{\mathcal{C},g,[a_{1},b_{1}]},\cdots,L_{\mathcal{C},g,[a_{k},b_{k}]}):{\mathcal{H}}(m^{O(g)},1^{2g-2-mO(g)})\to\mathbb{N}^{k}_{0}$
converges jointly in distribution to a vector of random variables with Poisson distributions of means $\lambda_{[a_{i},b_{i}]}$, where
$\lambda_{[a_{i},b_{i}]}=\frac{(m+1)\pi(b_{i}^{2}-a_{i}^{2})}{b^{\prime}b^{\prime\prime}}.$